OpenStudy (anonymous):

You invest an initial $100 in an account that has an annual interest rate of 3%, compounded quarterly. How much money will you have in the account after 20 years? Round your answer to the nearest whole number.

OpenStudy (whpalmer4):

Do you know the formula for compound interest?

OpenStudy (whpalmer4):

\[FV = PV(1+\frac{r}{n})^{nt}\]where \(FV\) is the future value, \(PV\) is the present value, \(r\) is the annual interest rate, expressed as a decimal, \(n\) is the number of compounding periods per year, and \(t\) is the number of years

OpenStudy (wolf1728):

Total = $100 * (1 + .03/4)^4*20 Total = $100 * 1.8180439804 Total = $181.80 and a calculator to check your answer is here: http://1728.org/compint.htm